# Are LLMs Naturally Good at Synthetic Tabular Data Generation?

### v2

@@ 4.4 RQ3: Can data generated by PAFT replace real data in downstream ML model training?
We next assesses effectiveness of the generated (i.e., synthetic) data by measuring how well discriminative models trained on synthetic data perform on their target discrimination task versus when trained on real data. As shown in Table 3, PAFT  is best or second best in over 80% of (dataset, method) combinations. Other research questions are discussed in the appendix, as indicated in Table 3.
5 Conclusion
@@
This work has brought LLMs closer to the goal of generating realistic synthetic datasets. By learning FDs and leveraging this information in the fine-tuning process, we are able to align the auto-regressive nature of LLMs with the ordering of columns necessary for generating quality synthetic data. While PAFT is quite broadly applicable by itself, it can be extended in several directions. First, what are other, perhaps more expressive, types of tabular constraints that can be utilized in the fine-tuning process? Second, what is the internal basis for regulating orders inside a transformer architecture and can we more directly harness it? Third, can we theoretically prove the (im)possibility of generating specific synthetic datasets by LLM architectures? Fourth, despite the numerous advantages of LLM in learning and generating tabular data, scalability remains an acknowledged challenge [11, 3, 14], encompassing concerns such as context window and training speed. And finally, privacy-preserving methods have been implemented in table generators based on GANs but remains understudied in LLM fine-tuning. These questions will be the focus of our future work.
@@
Appendix A Extended Methodology
A.1 Method Overview
Figure 6 shows an overview of the proposed permutation aided fine-tuning approach (PAFT ). A typical workflow is 1) Textual Encoding (Section 3.1) 2) Functional Dependency (FD) discovery (Section 3.2) 3) FD Distillation ( Section 3.2)and 4) The Feature Order Permutation Optimization ( Section 3.3). Fine-tuning and sampling Strategy is explained in Section 3.4.
Figure 6: Overview of the proposed Permutation-Aided Fine-tuning (PAFT ) approach.
A.2 Extended Algorithm
Algorithm 1 details FD distillation procedures corresponding to the text in Section 3.2 and Algorithm 2 details Feature Order Permutation Optimization procedures corresponding to the text in Section 3.3.
Algorithm 1 FD Distillation with Schema-Level FDs
1:List of Schema-level FDs, $\mathcal{S}$
2:Column Dependency Graph, $\mathcal{G}(\mathcal{V},\mathcal{E})$
3:$\mathcal{G}\leftarrow\emptyset$
4:for $fd\in\mathcal{S}$ do
5:     $LHS,RHS\leftarrow fd$
6:     if $LHS.length=1$ and $RHS.length=1$ then
7:         for $u\in LHS$ do
8:              for $v\in RHS$ do
9:                  $\mathcal{G}.\text{add\_edge}(u,v)$triangleright Case 1
10:     else if $LHS.length>1$ and $RHS.length=1$ then
11:         for $u\in LHS$ do
12:              $\mathcal{G}.\text{add\_edge}(v,u)$triangleright Case 2
13:     $\triangleright$ $RHS.length>1$, Case 3
14:         for $v\in RHS$ do
15:              if $LHS.length=1$ then
16:                  $\mathcal{G}.\text{add\_edge}(u,v)$triangleright Go to Case 1
17:              else
18:                  for $u\in LHS$ do
19:                       $\mathcal{G}.\text{add\_edge}(v,u)$triangleright Go to Case 2
Algorithm 2 Feature Order Permutation Optimization
1:Column Dependency Graphs $\mathcal{G}(\mathcal{V},\mathcal{E})$
2:Optimal Feature Order Permutation $\mathbf{k}$
3:if $|\mathcal{E}|$ = $\emptyset$ then
4:     Return $\mathbf{k}\leftarrow arbitrary\_permutation(\mathcal{V})$
5:if $\mathcal{G}$ is not a Directed Acyclic Graph (DAG) then $\triangleright$ Phase 1
6:     $\mathcal{G}\leftarrow strongly\_connected\_components(\mathcal{G})$
7:$\mathbf{k}\leftarrow topological\_ordering(\mathcal{G})$triangleright Phase 2
8:$\mathbf{k}\leftarrow arbitrary\_SCC\_expansion(\mathcal{G})$triangleright Phase 3
9:Return $\mathbf{k}$
Appendix B Additional Research Questions and Case Studies
B.1 RQ4: Does PAFT adhere to real distribution and possess mode diversity?
In this question, we evaluate how closely the density and diversity of the true data distribution are matched by PAFT generated data using correlation metrics and density-based distance metrics. Specifically, we employ Kolmogorov-Smirnov Test (KST) to evaluate the density estimate of numerical columns, and the Total Variation Distance (TVD) for categorical columns. The results for density estimate similarity are detailed in Table 4.
When calculating the correlation between columns, we employ Pearson correlation for numerical columns and contingency similarity for categorical columns. The results for correlation based analysis are detailed in Table 5. The above metrics can be calculated and normalized into the interval of 0-1 with the use of the sdv metric library [10], with 1 indicating that synthetic data perfectly matches real data distribution. Hence, evaluation score allows for easy comparison of data columns, even when they have different types and metrics.
Tables 4 and 5 demonstrate that the PAFT synthetic data closely matches the real data in terms of univariate distribution and bivariate correlation, outperforming the baseline. Visual examples are depicted in Figure 7. PAFT has the ability to generate a wide range of diversity, encompassing both continuous and discrete variables, which closely resembles real data.
Figure 7: Column distributions for each dataset generated by CTGAN, CopulaGAN, GReaT, and PAFT . Overall, PAFT has the closest distribution to real data compared to other synthesis methods. PAFT also showcase the ability to generate a wide range of diversity.
Dataset CTGAN CopulaGAN TabSyn GReaT PAFT
Adult 0.81$\pm$0.01 0.92$\pm$0.01 0.98$\pm$0.00 0.88$\pm$0.00 0.90$\pm$0.00
Beijing 0.89$\pm$0.01 0.79$\pm$0.01 0.98$\pm$0.00 0.93$\pm$0.00 0.97$\pm$0.00
California 0.87$\pm$0.03 0.77$\pm$0.01 0.98$\pm$0.00 0.89$\pm$0.00 0.83$\pm$0.00
US-locations 0.83$\pm$0.02 0.82$\pm$0.00 0.96$\pm$0.00 0.93$\pm$0.00 0.97$\pm$0.00
Seattle 0.83$\pm$0.01 0.73$\pm$0.02 0.93$\pm$0.00 0.90$\pm$0.00 0.93$\pm$0.00
Travel 0.84$\pm$0.01 0.90$\pm$0.02 0.93$\pm$0.01 0.93$\pm$0.01 0.93$\pm$0.01
Table 4: Error rate (%) of column-wise density estimation [10]. Bold Face represents the best score on each dataset. Higher values indicate more accurate estimation (superior results). PAFT outperforms the best generative baseline model in most case. The best results are marked in bold, the second-best results are underlined.
Dataset CTGAN CopulaGAN TabSyn GReaT PAFT
Adult 0.81$\pm$0.02 0.86$\pm$0.01 0.93$\pm$0.00 0.80$\pm$0.01 0.78$\pm$0.00
Beijing 0.92$\pm$0.01 0.94$\pm$0.01 0.99$\pm$0.00 0.95$\pm$0.00 0.98$\pm$0.01
California 0.84$\pm$0.00 0.87$\pm$0.01 0.97$\pm$0.00 0.87$\pm$0.01 0.91$\pm$0.02
US-locations 0.50$\pm$0.02 0.55$\pm$0.00 0.93$\pm$0.00 0.89$\pm$0.00 0.94$\pm$0.00
Seattle 0.74$\pm$0.02 0.72$\pm$0.01 0.80$\pm$0.01 0.76$\pm$0.03 0.81$\pm$0.01
Travel 0.77$\pm$0.02 0.80$\pm$0.02 0.87$\pm$0.01 0.85$\pm$0.01 0.82$\pm$0.05
Table 5: Error rate (%) of pair-wise column correlation score [10]. Bold Face represents the best score on each dataset. PAFT outperforms the best baseline model in most case. The best results are marked in bold, the second-best results are underlined.
B.2 RQ5: Does the synthetic data generated by PAFT pass the sniff test?
Similar to the analysis conducted in recent work [3] (GReaT), we employ the random forest (RF) algorithm to train discriminators to distinguish real data (labelled as True) and synthetically generated data (labeled as False). Subsequently, we test performance on an unseen set (consisting of 50% synthetically generated data and 50% real data). In this experiment, scores represent the percentage of correctly classified entities. In this case, an ideal accuracy score would be close to 50%, which means the discrimniator fails to distinguish between real and synthesized data. Scores are shown in Table 6 and indicate that the data generated by PAFT is most indistinguishable from real data, even by powerful discriminative models.
Method CTGAN CopulaGAN TabSyn GReaT PAFT
Beijing 99.16$\pm$ 0.08% 98.69$\pm$ 0.39% 50.97$\pm$0.0651.1$\pm$ 0.08% 50.09$\pm$ 0.05%
US-locations 99.94$\pm$ 0.03% 97.74$\pm$ 0.22% 51.97$\pm$0.1850.47$\pm$ 0.07% 50.01$\pm$ 0.01%
California 98.35$\pm$ 0.2% 86.64$\pm$ 0.67% 50.64$\pm$0.1553.74$\pm$ 0.27% 49.89$\pm$ 0.03%
Adult 94.43$\pm$ 0.53% 59.82$\pm$ 0.9% 51.64$\pm$0.1451.12$\pm$ 0.26% 48.75$\pm$ 0.03%
Seattle 87.61$\pm$ 1.06% 85.7$\pm$ 2.0% 50.12$\pm$0.968.27$\pm$ 1.34% 47.21$\pm$ 0.48%
Travel 77.96$\pm$ 1.4% 74.14$\pm$ 1.64% 50.66$\pm$1.9762.49$\pm$ 1.2% 48.18$\pm$ 0.81%
Table 6: Discriminator Performance (%): Comparison of synthesized data from CTGAN, CopulaGAN, TabSyn, GReaT, and PAFT . The scores stand for the accuracy for detecting real or fake data, where the ML models are trained using 50% real data and 50% random data. An ideal accuracy score is 50, indicating the model cannot distinguish between real and synthesized data. The best results are marked in bold, the second-best results are underlined.
B.3 RQ6: Can PAFT enhance the stability in generating high quality tables, resulting in a faster sampling phase?
When it comes to the comparison between fine-tuning approaches in LLM, the quality of the generated table rows can be reflected in the sampling process. This is particularly evident when generating an i.i.d row that involves auto-regressive generation. If the generated row cannot be decoded back into a real table row, then another sample needs to be redone. Given that the device condition and hyperparameters of the GReaT and PAFT are identical, a shorter sampling time indicates a higher probability of accepting a generated row, meaning improved generation quality. (See Figure 7.) Furthermore, the time measurement is based solely on a single-core GPU to ensure a fair comparison of different baselines. Using multiple GPU parallel computing can significantly speed up the fine-tuning process in practice.
Admittedly, PAFT and all LLM fine-tuning based table generators share the identical challenge of time consuming (both training and sampling), comparing to classic DL or GAN based table generators. In exchange for the investment of time, there are several advantages to consider. These include the elimination of data preprocessing which is a significant time cost for a human expert in charge of data preparation, a deep understanding of real-world knowledge, and finally DL-bsaed generators have been shown to fail when the table generation task involves generating columns with textual sentences as their values.
GReaT - Sampling Time PAFT  - Sampling Time
Adult 37:47 min 5:15 min
Beijing 5:24 min 5:21 min
California Housing 4:16 min 2:58 min
US-locations 58 sec 58 sec
Seattle 10 sec 11 sec
Travel 5 sec 5 sec
Table 7: For certain datasets that pose difficulties for auto-regressive generation (such as Adult), PAFT can significantly enhance the quality of the generation process, resulting in reduced generation time. For typical datasets with fewer challenges, the time-efficiency performance of GreaT and PAFT is comparable. (The standard variance of time in the five random experiments was smaller than one secondary unit, so it has been omitted).
B.4 Case Study: Influence of Statistical and Semantic Factors
Figure 8: Semantic Complexity: using the random order permutation to modeling the mixture of states data is more challenging when the rectangularity index (left) and compactness index (right) increase.
Figure 9: Statistical Complexity can be figured out by analyzing the distribution of the data. For instance, California Housing data is simpler to model concerning the Functional Dependency as it only includes the longitude and latitude for a single state.
Statistical Factor The occurrence of functional dependency in a table is influenced by various factors, including the number of rows and columns, the distinct values in the columns, the relationship between columns, and the presence of duplicate rows, etc. Figure 1 and 8 shows different dataset may have different level of statistic difficulties.
Semantic Factors The acquisition of semantic factors is typically not achievable from direct observation of the data’s appearance. Typically, this implies that the data’s worth will be influenced by real-world expertise in a specific field. For instance, map coordinates are influenced by the geopolitical borders of actual countries and states. Similarly, even if only a subset of data points from a mathematical function are observed, there is a need to comprehend the complete representation of that particular mathematical function.
Fig. 9 shows the difficulty of capturing the functional dependency can also leaded by the semantic conext of a sub-class in a mixture dataset, such as the state shape and geo location distribution. For this case, previous Fig. 5 have already shown the improvement of utilizing PAFT .
Figure 10: For a composite dataset, comparison of state-specific violation rates for different synthetic data generation approaches. Here, the states (x-axis) are sorted based on increasing violations. PAFT significantly mitigates state-specific violations in a composite dataset.
Appendix C Data and Experiment Description
C.1 Experimental Setup
Dataset. We evaluated the efficiency of PAFT through experiments on six real datasets commonly used in synthetic table generation studies (GReaT, CTGAN, etc.), as well as a set of four simulated datasets: Beijing [7], US-locations [13], California Housing [24], Adult Income [2], Seattle [34], and Travel [37], Simulated (Algorithm 3). These real-world datasets come from diverse domains and vary in size. The range of the number of functional dependencies spans from 0, indicating complete independence between columns, to around 400, indicating a high level of interdependence. These data also demonstrate the diverse combinations of categories and numerical columns. The simulated data is customized to adhere to the given functional dependence schema, thus explicitly emphasizing the degree to which a model can precisely represent the functional relationship. The simulated data has four distinct versions, denoted by the variable $k$, which represents the unique values in the $d$ column. As the value of $k$ increases, the data becomes more complex, which has been demonstrated to make training the generative model more challenging, as evidenced by all the experimental results presented in this paper. In particular, the functional dependency graph of the simulated data is [$a\rightarrow b$, $a\rightarrow c$,$b\rightarrow c$,$b\rightarrow d$,$a\rightarrow d$,$b\rightarrow d$]. Table 8 provides the FD characteristics of each dataset, while Table 9 provides an more detailed overview.
Training and Testing. To prevent any data leakage, we partitioned the data sets into 80% training sets and 20% test sets. All models are trained or fine-tuned on the same training data samples. All models undergo cross-validation using 5 generated data sets to validate their results. One advantage of using LLM for creating tabular data is that there is no need for any complex data preparation. This means that the feature names and values are used just as they are supplied in the original data sets.
Dataset Beijing US-locations California Housing Adult Income Seattle Travel Simiulated
# of FDs 157 7 362 78 10 0 6
Table 8: The number of functional dependencies identified using automated FD discovery (Sec 3.2).
Dataset #Rows #Cat #Num #FD
Beijing 43,824 1 12 157
US-locations 20,400 3 2 7
California 20,640 8 0 362
Adult 32,561 9 6 78
Seattle 2,016 2 6 10
Travel 954 4 3 0
Simulated,k=[1,5,10,15] 10,000 4 0 6
Table 9: Dataset Descriptors (number of rows, categorical columns, numerical columns, and FDs).
Algorithm 3 Building simulated data: Given a dependency graph G, setting values for a table with $n$ rows and $m$ columns with different statistic complexity based on the initial unique value $k$ for the complexity in the root column of the functional dependency chain.
1:procedure SetValues($n,m,G,k$)
2:     $table[n][m]\leftarrow\{\}$
[…]
4:     $unique\_value\leftarrow\{\}$
5:     for $vertex\in topo\_order$ do
6:         $top\_value[vertex]=2^{max\_layer-node\_layer[vertex]}$
7:     for $j\leftarrow 1$ to $m$ do
8:         for $i\leftarrow 1$ to $n$ do
9:              $table[i][j]\leftarrow x_{i\%(top\_value[j]*k)}$
Baselines. In benchmarking suite, we have baselines that consist of current deep learning approaches for synthetic data generation (CTGAN [40], CopulaGAN [40]), and the most advanced LLM fine-tuning synthetic table generator GReaT [3]. To guarantee an equitable comparison, we employ the Distill-GReaT model for both techniques in all tests, and adjust the hyperparameters as advised by the official GitHub website of GReaT. It ought to mention that GReaT utilizes textual encodings with random feature order permutations. This implies that each sample will undergo a different random order during the training and sampling process. This strategy appears similar to the edge case in Algorithm 2, but in fact, they are distinct. When the FD set is empty, PAFT will suggest a random permutation. Nevertheless, this permutation serves as a comprehensive guide for all samples after an in-depth assessment of FD.
C.2 Reproducibility detail
Baselines. Each baseline (CTGAN, CopulaGAN, TabSyn, GReaT) sticks to the recommended hyperparameters and utilizes officially released API tools: Synthetic Data Vault [28] and GReaT [3]. As for the fair comparison of GReaT and PAFT, the LoRA fine-tuning parameters are set as: Lora attention dimension $r=16$, alpha parameter for Lora scaling $lora\_alpha=32$, The names of the modules to apply the adapter to $target\_modules=c\_attn$, The dropout probability for Lora layers $lora\_dropout=0.05$, $bias=none$.
Computational Resources. To ensure fairness in the comparison between the baselines, all baseline models and experiments were executed on a single Tesla P100-PCIE-16GB GPU.
Parameters for MLE and Discriminator Models. We utilize neural network, linear/logistic regression, and random forest models from the Scikit-Learn package for the ML efficiency and discriminator experiments. The exact hyperparameters for each model are detailed in Table 10. Every result is evaluated through the process of 5-fold cross-validation.
RF LR NN
n_est max_depth max_iter max_iter hidden_layer_sizes learning_rate
Classification 100 None 100 300 (150, 100, 50) 0.001
Regression 100 None 100 300 (150, 100, 50) 0.001
Table 10: The parameters we used for MLE and Discriminator models remain the same across all datasets.
Appendix D Related Work
Owing to the ubiquity of tabular data, the synthetic generation of this type of data in traditional machine learning research. Various approaches have been developed for tabular data generation.
Tabular Data Generation with Neural Networks (i.i.d. rows). Lei et al. [40] proposed CTGAN where rows are independent of each other; a conditional GAN architecture ensures that the dependency between columns is learned. Tabsyn [43] also generates independent rows but with a diffusion approach.
Tabular Data Generation with Neural Networks (non i.i.d. rows). DoppelGanger [19] uses a combination of an RNN and a GAN to incorporate temporal dependencies across rows but this method has been tested in traditional, low-volume settings such as Wikipedia daily visit counts. For high-volume applications, STAN [42] utilizes a combination of a CNN and Gaussian mixture neural networks to generate synthetic network traffic data. GraphDF  [6] conducts multi-dimensional time series forecasting. GOGGLE  [20] employs a generative modeling method for tabular data by learning relational structures.
Use of Language Models (LLMs) for tabular data generation. Most modern LLMs are based on the transformer architecture [39] with parameters ranging from few millions to billions [15], and researchers have developed creative ways to harness LLMs in traditional machine learning and data contexts. LIFT [11] initially transforms a table row into a sentence, such as ‘An Iris plant with sepal length 5.1cm, sepal width 3.5cm’, and employs an LLM as a learning model for table classification, regression, and generation tasks. GReaT [3], introduced earlier, utilizes a GPT-2 model that has been fine-tuned using a specific corpus for synthetic data generation. A general benefit of utilizing LLMs is the elimination of customized preprocessing pipelines.
Feature Ordering. Although not well-studied in the context of tabular data generation, the notion of feature ordering has been investigated in the context of graph-to-text translation [36] wherein to learn effective graph encodings, vertices are linearized via combinations of different graph traversal mechanisms, e.g., topological & breath-first strategies [12], and top-down & bottom-up approaches [32]. As a second example, ‘permutation-invariant tabular data synthesis [47] examines the influence of the arrangement of table columns on convolutional neural network (CNN) training and organizes them according to the correlation among columns. Nevertheless, it is important to acknowledge that relying on mere correlation to establish column orders can be limiting. There are also several approaches (e.g., [17, 9]) that synthesize, discover, or aggregate features from relational databases, leveraging order information when possible, for use in machine learning pipelines. It worth to note that even in the LLM community, the task of context sorting for LLM prompting is not trivial and has gained significant attention lately [8].
Mining and Modeling Functional Dependencies. Yunjia et al. [44] relax the notion of strict functional dependencies to include noisy functional relationships by utilizing probabilistic graphical models. Chen et al. [5], in their FakeTables approach use the discovery of functional dependencies in a GAN formulation; they first use a generator to create a set of columns (set A) and an autoencoder to cast another set (set B), which are then used by a discriminator to calculate the gradient loss. Muralidhar et al. [23] proposed to use Granger causality to incorporate functional invariants across multiple time series. The area of FD discovery and of data mining with tabular data have both been extensively studied [21, 29, 46].

### v3

@@ 5.4. RQ4: Can data generated by PAFT replace real data in downstream ML model training?
We next assess the effectiveness of the generated (synthetic) data by comparing the performance of discriminative models trained on synthetic data versus real data for their target tasks. Models tested include random forests (RF), linear regression (LR), and multi-layer perceptron (NN). As shown in Table 5, PAFT  is best or second best in over 80% of (dataset, method) combinations.
5.5. RQ5: Do the data sets generated by PAFT adhere to real distributions and possess mode diversity?
We also evaluate how closely the density and diversity of the true data distribution are matched by PAFT generated data using correlation metrics and density-based distance metrics. Specifically, we employ the Kolmogorov-Smirnov Test (KST) to evaluate the density estimate of numerical columns, and the Total Variation Distance (TVD) for categorical columns. When calculating the correlation between columns, we employ Pearson correlation for numerical columns and contingency similarity for categorical columns. The results for both density estimate similarity and correlation based analysis are detailed in Table 6. As shown, PAFT synthetic data closely matches the real data in terms of univariate distribution and bivariate correlation, outperforming the baseline.
Table 6. Low-order statistics (Zhang et al., 2023b) of column-wise data density and pair-wise column correlation2. Scores range from 0 to 1. Higher values indicate more accurate estimation. PAFT outperforms the best generative baseline model in most case. The best results are marked in bold, the second-best results are underlined. Results are averaged over five random runs.
Single-Column Shape Score ($\uparrow$)
Dataset CTGAN Co.GAN TabSyn GReaT PAFT
Adult 0.81 0.92 0.98 0.88 0.90
Beijing 0.89 0.79 0.98 0.93 0.97
California 0.87 0.77 0.98 0.89 0.83
US-locations 0.83 0.82 0.96 0.93 0.97
Seattle 0.83 0.73 0.93 0.90 0.93
Travel 0.84 0.90 0.93 0.93 0.93
Two-Column Pair Trends score ($\uparrow$)
Dataset CTGAN Co.GAN TabSyn GReaT PAFT
Adult 0.81 0.86 0.93 0.80 0.78
Beijing 0.92 0.94 0.99 0.95 0.98
California 0.84 0.87 0.97 0.87 0.91
US-locations 0.50 0.55 0.93 0.89 0.94
Seattle 0.74 0.72 0.80 0.76 0.81
Travel 0.77 0.80 0.87 0.85 0.82
Visual examples are depicted in Figure 4. PAFT has the ability to generate a wide range of diversity, encompassing both continuous and discrete variables, which closely resembles real data.
Figure 4. Column distributions visualization for each dataset generated by CTGAN, CopulaGAN, GReaT, and PAFT . The top row displays examples of numerical columns, while the bottom row presents examples of categorical columns. Overall, PAFT (Blue) has the closest distribution to real data (Red) compared to other synthesis methods. PAFT also showcase the ability to generate a wide range of diversity.
5.6. RQ6: Do newer generations of LLMs obviate the need for PAFT ?
The choice of foundation models for synthetic table generation involves two key considerations: in-context learning and fine-tuning, especially with large-parameter models such as GPT-4 and LLaMA. A recent survey (Fang et al., 2024) indicates that the in-context learning approach of GPT-4 is well-suited for augmenting tabular data in low-data regimes, as highlighted by CLLM (Seedat et al., 2024). However, GPT-4 suffers from limitations such as the disappearance of column-wise tail distributions and a low success rate in accurately extracting output cell values (Fang et al., 2024).
In contrast, the state-of-the-art GReaT (Borisov et al., 2022) and other synthetic data generation approaches (Solatorio and Dupriez, 2023; Zhang et al., 2023a; Zhao et al., 2023) typically employ smaller models like GPT-2 or DistilGPT2, which effectively address attribute encoding while reducing feature engineering efforts. These smaller models already outperform traditional GAN and traditional statistics-based approaches. For example, DistilGPT2 can be trained using LoRA on GPUs as modest as the Tesla P100.
Importantly, fine-tuning larger models such as GPT-4 entails significant computational cost. For instance, fine-tuning a GPT-4 model for three epochs on a dataset with 50k rows and a five-column table costs approximately \$300 (OpenAI) or requires equivalent high-end GPU resources. Therefore, while the same fine-tuning schema can be applied to models like GPT-4 or LLaMA, it is not a cost-effective solution compared to the solutions considered here.
In summary, newer generation LLMs do not obviate the need for PAFT . Our method directly addresses the impedance mismatch between autoregressive LLMs and synthetic table generation—particularly by preserving functional dependencies through permutation-aware fine-tuning—a challenge that persists even with advanced models.
6. Conclusion
@@
This work has brought LLMs closer to the goal of generating realistic synthetic datasets. By learning FDs and leveraging this information in the fine-tuning process, we are able to align the auto-regressive nature of LLMs with the ordering of columns necessary for generating quality synthetic data. While PAFT is quite broadly applicable by itself, it can be extended in several directions. First, what are other, perhaps more expressive, types of tabular constraints that can be utilized in the fine-tuning process? Second, what is the internal basis for regulating orders inside a transformer architecture and can we more directly harness it? Third, can we theoretically prove the (im)possibility of generating specific synthetic datasets by LLM architectures? Fourth, despite the numerous advantages of LLM in learning and generating tabular data, scalability remains an acknowledged challenge (Dinh et al., 2022; Borisov et al., 2022; Fang et al., 2024), encompassing concerns such as context window and training speed. And finally, privacy-preserving methods have been implemented in table generators based on GANs but remains understudied in LLM fine-tuning. These questions will be the focus of our future work.